OpenStudy (anonymous):

(a+b)(a2-ab+b2)=a3+b3 Why is that equal?

OpenStudy (shadowlegendx):

Is that a times 2 or a to the second power?

OpenStudy (anonymous):

it wasn't time 2

OpenStudy (shadowlegendx):

\[(a + b)(a ^{2} - ab + b ^{2} = a ^{3} + b ^{3}\]

OpenStudy (shadowlegendx):

Looks like that?

OpenStudy (anonymous):

it was just like that. Is that possible? (a+b)(a2-ab+b2)=a3+b3

OpenStudy (shadowlegendx):

Why? Is it asking you if it's true or not?

OpenStudy (anonymous):

(a+b)(a2-ab+b2)=a3+b3

OpenStudy (unklerhaukus):

(a + b)(a^2 - ab + b^2) = a(a^2 - ab + b^2) + b(a^2 - ab + b^2) = a^3-a^2b+ab^2 + a^2b -ab^2 +b^3 = a^3 + b^3